FIND THE GCF OF 6 18 AND 30
Answer by funmath(2873)

FIND THE GCF OF 6 18 AND 30
6 is the GCF, because it's the biggest factor of all three numbers
6=6*1
18=6*3
30=6*5
Happy Calculating!!!